A bag of 31 tulip bulbs contains 10 red tulip? bulbs, 12 yellow tulip? bulbs, and 9 purple tulip bulbs.

?(a) What is the probability that two randomly selected tulip bulbs are both? red?

?(b) What is the probability that the first bulb selected is red and the second? yellow?

?(c) What is the probability that the first bulb selected is yellow and the second? red?

?(d) What is the probability that one bulb is red and the other? yellow?
Answer #1

a)

probability that two randomly selected tulip bulbs are both? red=P(first red)*P(second red|first red)=(10/31)*(9/30)=3/31

=0.0968

b)probability that the first bulb selected is red and the second? yellow=(10/31)*(12/30)=4/31=0.1290

c)

probability that the first bulb selected is yellow and the second? red=(12/31)*(10/30)=4/31=0.1290

d)

probability that one bulb is red and the other? yellow =P(first yellow and second red)+P(irst red and second yellow)

=(4/31)+(4/31)=8/31=0.2581